OpenStudy (anonymous):

Each of the letters in the word MATHEMATICS is represented by a letter tile in a bag. Stacey picks three letters, without replacement. To the nearest whole percent, which is the probability that she'll get all vowels? answer choices: 2% -3%-5%-6%

OpenStudy (anonymous):

yolo

OpenStudy (poopsiedoodle):

Really? You only need to use logic on this one.

OpenStudy (poopsiedoodle):

This is a joke, right?

OpenStudy (poopsiedoodle):

seems like someone may be trying to cheat...

OpenStudy (kropot72):

There are 4C3 ways of drawing 3 vowels out of the 4 vowels that are available. There are 11C3 ways of drawing 3 letter tiles from the 11 tiles in the bag. The required probability is given by \[\frac{4C3}{11C3}=4\times \frac{3!8!}{11!}=you\ can\ calculate\]